OpenStudy (agent47):

Quick probability question, just want to check my approach, please don't post the answer.

OpenStudy (agent47):

Supose you have 5 bins and 5 balls. Each of the balls has probability of 1/5 of landing in the bin when thrown. So, suppose you try to throw all five balls into the bins. What is the probability that at least one of the bins is empty? So, what I'm doing is: P(at least 1 empty) = 1 - P(none empty) P(none empty, meaning a ball in each of the bins) = 1/5*5/5 * 1/5*4/5 * 1/5*3/5 * 1/5*2/5 * 1/5*1/5

OpenStudy (agent47):

Does this look right?

OpenStudy (inkyvoyd):

I got 5/5*4/5*3/5*2/5*1/5

OpenStudy (inkyvoyd):

well 1 minus that

OpenStudy (agent47):

Guys, but there's a probability of 1/5 of a ball even landing in the bin

OpenStudy (inkyvoyd):

P(none empty)=P(all full)

OpenStudy (inkyvoyd):

oh, they aren't ordered lol

OpenStudy (agent47):

Interpreting the problem is a pain in the bum... hang on ill attach a screenshot of the actual problem

OpenStudy (inkyvoyd):

I think I"m right about it though, 5/5 chances to get into a empty, then 4/5 chances to get into an empty, then 3/5 and so on and so forth

OpenStudy (inkyvoyd):

so it's 1-5!/5^5

OpenStudy (agent47):

I mean disregard the numbers, I just changed them around a bit, in order to not be accused of cheating on my hw - but yea it's the question in the screenshot.

OpenStudy (inkyvoyd):

ya, but same logic... try PA, the mod there will help

OpenStudy (agent47):

PK halp, I need a third opinion.

Parth (parthkohli):

That "1/4 probability" sentence could also mean that a ball is equally likely to fall in any of the boxes - no biases, that is.

Parth (parthkohli):

Like when we do problems about fair coins, they may specify something like the chance of landing a head is 1/2.

OpenStudy (agent47):

Yea, I mean in that case it would have to be 1/4!

OpenStudy (agent47):

Half the time I have trouble interpreting the problem

Parth (parthkohli):

I know, it's a pain in the posterior. Plus it uses the phrase "any particular box", so it does seem like this means nothing but that a ball is equally likely to fall in all four boxes.

OpenStudy (inkyvoyd):

yeah that's 80% of probability

Parth (parthkohli):

So Inky's answer seems fair enough to me.

OpenStudy (agent47):

Okay fine that explains it - I originally took that approach, but then thought I had to do something about that 1/4 probability.